OpenStudy (anonymous):

six minus 1 over four ?? HELPP

OpenStudy (anonymous):

6-1/4 this is the equation

OpenStudy (anonymous):

6 - 1/4 = 5 3/4

OpenStudy (anonymous):

if the 6-1 is in prencthesis() then it is 5/4

OpenStudy (anonymous):

if it is not, then it is 6-1/4=23/4; if you still have questio ask me :)

OpenStudy (anonymous):

(6/1) - (1/4) (24/4) - (1/4) = 23/4